# On the optimal prediction of extreme events in heavy-tailed time series with applications to solar flare forecasting

### v1

@@ A.2.7 Half-Spaces in $\mathbb{R}^{d}$
$\geq\|\varphi^{(k)}-\varphi\|^{-1/2},$
@@ A.3 Additional Information for Section 4
A.3.1 Autoregressive Model Simulation Details
For the simulation study in Section 4.1, data was generated from the AR(5) model
with the $\epsilon_{t}$’s being iid $t$-distributed with one degree of freedom. The corresponding AR polynomial has roots $-3\pm 1i$, $4\pm 2i$, and $2$; since these roots lie outside the closed unit disk, (A.43) has a unique solution that is stationary and causal. Set
$\phi=(0.3,0.19,-0.035,-0.01,0.0025).$
Next, one hundred times, a time series of length $10^{4}+10^{6}$ was generated from the model in (A.43); the first $10^{4}$ observations were reserved for training, and the last $10^{6}$ were reserved for testing. Inspired by Davis et al. (1992), LAD estimation was used on the training set to estimate $\widehat{\phi}$ of $\phi$ for the non-oracle predictors; the order of the model was assumed to be known. Essentially, a median regression was carried out in which the response was the observation one step ahead and the covariates were the five most recent observations. Then, for the three predictors described in Section 4.1, one step-ahead prediction was performed for each $p$; prediction worked the same way for each $p$.
$\phi^{\top}Y_{t:(t-4)}>F_{\widehat{Y}_{t+h}(\phi)}^{\leftarrow}(p).$
$\widehat{\phi}^{\top}Y_{t:(t-4)}>\widehat{F}_{\widehat{Y}_{t+h}(\widehat{\phi})}^{\leftarrow}(p)),$
where $\widehat{F}_{\widehat{Y}_{t+h}(\widehat{\phi})}^{\leftarrow}(p))$ was computed as the $p$th sample quantile of all linear combinations of the form $\widehat{\phi}^{\top}Y_{s:(s-4)}$ that could be computed from the training set.
$\widehat{\phi}^{\top}Y_{t:(t-4)}>\widetilde{F}_{\widehat{Y}_{t+h}(\widehat{\phi})}^{\leftarrow}(p)),$
where $\widetilde{F}_{\widehat{Y}_{t+h}(\widehat{\phi})}^{\leftarrow}(p))$ was computed from the training set using the following procedure. For each of ten equally-spaced values between the third quartile of the training linear combinations and the tenth largest linear combination, a GP (Generalized Pareto) model was fit to the declustered excesses above the value. The excesses were the differences between the training observations above the value and the value, and they were declustered using the intervals method introduced in Ferro & Segers (2003). The purpose of declustering was to extract excesses that could reasonably be viewed as independent, which was important since the GP model was fit using maximum likelihood estimation. The GP distribution with scale parameter $\sigma>0$ and shape parameter $\xi$ has a distribution function that equals
$G(y)=1-\left(1+\frac{\xi y}{\sigma}\right)^{-1/\xi}$
for $y>0$ and zero otherwise; GP distributions are frequently used to model excesses above a threshold (Coles 2001).
$\widetilde{F}_{\widehat{Y}_{t+h}(\widehat{\phi})}^{\leftarrow}(p))=\tau+\widehat{G}^{\leftarrow}\left(\frac{p-p_{0}}{1-p_{0}}\right),$
Once all predictions were made for the test set, the precision was computed as the number of exceedances that were predicted to be exceedances divided by the number of predicted exceedances. Since there were 100 training set-test set pairs, there were 100 precisions for each pair of a predictor and a value of $p$. These precisions are displayed in boxplots in Figure 4.1. The dashed orange line represents the asymptotic precision of the optimal predictor, which was computed by deriving the causal representation of the solution to (A.43) and then applying Theorem 3.4 using the first $10^{6}$ terms in each series.
A.3.2 An Expanded Table of Data Analysis Results
@@
$h$ $p$ Baseline FARIMA AR(168) Baseline FARIMA AR(168)
1 0.90 0.492 0.448 0.354 0.435 0.392 0.276
1 0.95 0.379 0.368 0.258 0.334 0.320 0.225
1 0.99 0.244 0.238 0.104 0.262 0.262 0.124
6 0.90 0.308 0.339 0.284 0.256 0.298 0.215
6 0.95 0.189 0.213 0.150 0.173 0.198 0.138
6 0.99 0.049 0.119 0.125 0.040 0.115 0.162
12 0.90 0.284 0.292 0.253 0.197 0.205 0.152
12 0.95 0.194 0.176 0.127 0.148 0.124 0.076
12 0.99 0.098 0.073 0.049 0.085 0.061 0.056
18 0.90 0.270 0.309 0.253 0.210 0.256 0.183
18 0.95 0.175 0.197 0.142 0.154 0.173 0.117
18 0.99 0.049 0.116 0.067 0.040 0.115 0.064
Table A.2: A summary of the performance of various predictors. For each pair of $h$ and $p$, the goal was to predict exceedance of the $p$th marginal quantile $h$ steps ahead. In each row, for each metric, the largest value is bolded.

### v2

@@ A.2.7 Regular Variation
Define $g:(0,1)\to\mathbb{R}$ by $g(p)=\mathbb{P}(X>F_{X}^{\leftarrow}(p),Y>F_{Y}^{\leftarrow}(p))$; note that $g$ is increasing. Let $\{p_{k}\}$ be a sequence in $(0,1)$ such that $p_{k}\uparrow 1$ as $k\to\infty$. For each $k$, set $n_{k}=\lfloor\frac{1}{1-p_{k}}\rfloor$; as $k\to\infty$, $n_{k}\to\infty$. We have
$\lambda=\lim_{k\to\infty}n_{k}\mathbb{P}(X>F_{X}^{\leftarrow}(1-1/n_{k}),Y>F_{Y}^{\leftarrow}(1-1/n_{k}))=\lim_{k\to\infty}\frac{g\left(1-1/n_{k}\right)}{1/n_{k}}.$ (A.47)
We also have
$n_{k}\leq\frac{1}{1-p_{k}}<n_{k}+1$
$\frac{1}{n_{k}+1}<1-p_{k}\leq\frac{1}{n_{k}},$ (A.48)
which implies that $1-p_{k}\sim\frac{1}{n_{k}}$ and $1-p_{k}\sim\frac{1}{n_{k}+1}$ as $k\to\infty$. It also follows from (A.48) that
$1-\frac{1}{n_{k}}\leq p_{k}<1-\frac{1}{n_{k}+1}$
$\frac{g\left(1-\frac{1}{n_{k}}\right)}{1-p_{k}}\leq\frac{g(p_{k})}{1-p_{k}}\leq\frac{g\left(1-\frac{1}{n_{k}+1}\right)}{1-p_{k}}$
$\frac{g\left(1-\frac{1}{n_{k}}\right)}{\frac{1}{n_{k}}}\cdot\frac{\frac{1}{n_{k}}}{1-p_{k}}\leq\frac{g(p_{k})}{1-p_{k}}\leq\frac{g\left(1-\frac{1}{n_{k}+1}\right)}{\frac{1}{n_{k}+1}}\cdot\frac{\frac{1}{n_{k}+1}}{1-p_{k}}.$
Letting $k\to\infty$ on the last line and applying (A.47) yields
$\lim_{k\to\infty}\frac{g(p_{k})}{1-p_{k}}=\lambda.$ (A.49)
Hence,
$\lambda(X,Y)$ $=\lim_{k\to\infty}\mathbb{P}(X>F_{X}^{\leftarrow}(p_{k})\mid Y>F_{Y}^{\leftarrow}(p_{k}))$
$=\lim_{k\to\infty}\frac{\mathbb{P}(X>F_{X}^{\leftarrow}(p_{k}),Y>F_{Y}^{\leftarrow}(p_{k}))}{\mathbb{P}(Y>F_{Y}^{\leftarrow}(p_{k}))}$
$=\lim_{k\to\infty}\left[\frac{g(p_{k})}{1-p_{k}}\cdot\frac{1-p_{k}}{\mathbb{P}(Y>F_{Y}^{\leftarrow}(p_{k}))}\right]$
$=\lambda,$
where the last line follows from (A.49) and Lemma A.5. ∎
We now prove Lemma 3.4.
Proof of Lemma 3.4.
Let
Set $\delta=\delta_{a}\vee\delta_{b}$. Since $\delta/2\leq 1$,
It follows that
where $\nu_{\epsilon}$ is defined by
$\nu_{\epsilon}(dt)=\kappa_{+}(t)\alpha|t|^{-\alpha-1}\,dt.$
$\lambda(X,Y)=\lim_{n\to\infty}n\mathbb{P}(X>x_{0}c_{n},Y>y_{0}c_{n})=\nu_{Z}((x_{0},\infty)\times(y_{0},\infty))$
for some $x_{0},y_{0}>0$ such that $x_{0}c_{n}\sim F_{X}^{\leftarrow}(1-1/n)$ and $y_{0}c_{n}\sim F_{Y}^{\leftarrow}(1-1/n)$ as $n\to\infty$. Examining the geometry of the support of $\nu_{A_{j}\epsilon_{j}}$ and considering cases will allow us to compute $\nu_{A_{j}\epsilon_{j}}((x_{0},\infty)\times(y_{0},\infty))$, which will then make it possible to compute $\nu_{Z}((x_{0},\infty)\times(y_{0},\infty))=\sum_{j=0}^{\infty}\nu_{A_{j}\epsilon_{j}}((x_{0},\infty)\times(y_{0},\infty))$.
Since $X$ is regularly varying and $p_{X}=p_{\xi(a)}>0$, $\overline{F}_{X}(F_{X}^{\leftarrow}(p))\sim 1-p$ as $p\uparrow 1$ by Lemma A.5. This implies that $\lim_{n\to\infty}n\mathbb{P}(X>F_{X}^{\leftarrow}(1-1/n))=1$. Suppose we could find an $x_{0}>0$ such that $\lim_{n\to\infty}n\mathbb{P}(X>x_{0}c_{n})=1$. It would then follow from Lemma A.6 that as $n\to\infty$, $F_{X}^{\leftarrow}(1-1/n)\sim x_{0}c_{n}$.
For any $x_{0}>0$,
$=\nu_{Z}((x_{0},\infty)\times\mathbb{R})$
$=x_{0}^{-\alpha}\nu_{Z}((1,\infty)\times\mathbb{R})$
We have
[…]
If $a_{j}<0$, then we need $t<1/a_{j}$, so
[…]
Hence, whether $a_{j}$ is zero, positive, or negative,
Using (A.50), in order to have $\lim_{n\to\infty}n\mathbb{P}(X>x_{0}c_{n})=1$, we must have
$1=x_{0}^{-\alpha}\sum_{j=0}^{\infty}\nu_{A_{j}\epsilon_{j}}((1,\infty)\times\mathbb{R})=x_{0}^{-\alpha}\sum_{j=0}^{\infty}\kappa_{+}(a_{j})|a_{j}|^{\alpha}=x_{0}^{-\alpha}\eta_{+}({a},{0}).$
Since $p_{X}=p_{\xi(a)}>0$, Lemma 3.3 implies that $\eta_{+}({a},{0})>0$. Thus, we can set $x_{0}=\eta_{+}({a},{0})^{1/\alpha}>0$. We have thus found an $x_{0}>0$ such that as $n\to\infty$, $x_{0}c_{n}\sim F_{X}^{\leftarrow}(1-1/n)$. In a similar way, we can show that when $y_{0}=\eta_{+}({b},{0})^{1/\alpha}$, then $y_{0}c_{n}\sim F_{Y}^{\leftarrow}(1-1/n)$ as $n\to\infty$.
Now,
$\lim_{n\to\infty}n\mathbb{P}(X>x_{0}c_{n},Y>y_{0}c_{n})$ $=\lim_{n\to\infty}n\mathbb{P}(c_{n}^{-1}Z\in(x_{0},\infty)\times(y_{0},\infty))$
$=\nu_{Z}((x_{0},\infty)\times(y_{0},\infty))$ (A.51)
$=\sum_{j=0}^{\infty}\nu_{A_{j}\epsilon_{j}}\left([(x_{0},\infty)\times(y_{0},\infty)]\cap{\text{span}}\{A_{j}\}\right).$ (A.52)
[…]
$=p_{\epsilon}\left(\frac{a_{j}^{\alpha}}{x_{0}^{\alpha}}\bigwedge\frac{b_{j}^{\alpha}}{y_{0}^{\alpha}}\right)$
Similarly, if $a_{j},b_{j}<0$, then
$\nu_{A_{j}\epsilon_{j}}([(x_{0},\infty)\times(y_{0},\infty)]\cap{\text{span}}\{A_{j}\})=\kappa_{+,+}(a_{j},b_{j})\left(\frac{|a_{j}|^{\alpha}}{\eta_{+}({a},{0})}\bigwedge\frac{|b_{j}|^{\alpha}}{\eta_{+}({b},{0})}\right).$ (A.53)
Substituting the expression on the right side of (A.53) in (A.52) yields
$\lim_{n\to\infty}n\mathbb{P}(X>x_{0}c_{n},Y>y_{0}c_{n})=\sum_{j=0}^{\infty}\kappa_{+,+}(a_{j},b_{j})\left(\frac{|a_{j}|^{\alpha}}{\eta_{+}({a},{0})}\bigwedge\frac{|b_{j}|^{\alpha}}{\eta_{+}({b},{0})}\right);$ (A.54)
the expression on the right side of (A.54) is the expression for $\lambda(\xi(a),\xi(b))$ in the statement of the lemma.
We next show that
$\lim_{n\to\infty}n\mathbb{P}(X>F_{X}^{\leftarrow}(1-1/n),Y>F_{Y}^{\leftarrow}(1-1/n))=\lim_{n\to\infty}n\mathbb{P}(X>x_{0}c_{n},Y>y_{0}c_{n}).$ (A.55)
Let $\epsilon>0$. Since $x_{0}c_{n}\sim F_{X}^{\leftarrow}(1-1/n)$ and $y_{0}c_{n}\sim F_{Y}^{\leftarrow}(1-1/n)$ as $n\to\infty$, for sufficiently large $n$,
$\frac{F_{X}^{\leftarrow}(1-1/n)}{x_{0}c_{n}},\frac{F_{Y}^{\leftarrow}(1-1/n)}{y_{0}c_{n}}\in(1-\epsilon,1+\epsilon).$
so
$F_{X}^{\leftarrow}(1-1/n)\in((1-\epsilon)x_{0}c_{n},(1+\epsilon)x_{0}c_{n})\ \text{and}\ F_{Y}^{\leftarrow}(1-1/n)\in((1-\epsilon)y_{0}c_{n},(1+\epsilon)y_{0}c_{n}).$
Thus, for large enough $n$,
$n\mathbb{P}(X>F_{X}^{\leftarrow}(1-1/n),Y>F_{Y}^{\leftarrow}(1-1/n))\geq n\mathbb{P}(X>(1+\epsilon)x_{0}c_{n},Y>(1+\epsilon)y_{0}c_{n}),$
which implies that
$\liminf_{n\to\infty}n\mathbb{P}(X>F_{X}^{\leftarrow}(1-1/n),Y>F_{Y}^{\leftarrow}(1-1/n))$ $\geq\liminf_{n\to\infty}n\mathbb{P}(X>(1+\epsilon)x_{0}c_{n},Y>(1+\epsilon)y_{0}c_{n})$
$\geq\liminf_{n\to\infty}n\mathbb{P}(c_{n}^{-1}Z\in(1+\epsilon)[(x_{0},\infty)\times(y_{0},\infty)])$
$\geq\nu_{Z}((1+\epsilon)[(x_{0},\infty)\times(y_{0},\infty)])$
$\geq(1+\epsilon)^{-\alpha}\nu_{Z}((x_{0},\infty)\times(y_{0},\infty)).$ (A.56)
In a similar way, we can show that
$\limsup_{n\to\infty}n\mathbb{P}(X>F_{X}^{\leftarrow}(1-1/n),Y>F_{Y}^{\leftarrow}(1-1/n))\leq(1-\epsilon)^{-\alpha}\nu_{Z}((x_{0},\infty)\times(y_{0},\infty)).$ (A.57)
Chaining together (A.56) and (A.57) and letting $\epsilon\downarrow 0$ yields
$\lim_{n\to\infty}n\mathbb{P}(X>F_{X}^{\leftarrow}(1-1/n),Y>F_{Y}^{\leftarrow}(1-1/n))=\nu_{Z}((x_{0},\infty)\times(y_{0},\infty)),$
which implies (A.55), because of (A.51). Combining (A.54) and (A.55), we get
$\lim_{n\to\infty}n\mathbb{P}(X>F_{X}^{\leftarrow}(1-1/n),Y>F_{Y}^{\leftarrow}(1-1/n))=\sum_{j=0}^{\infty}\kappa_{+,+}(a_{j},b_{j})\left(\frac{|a_{j}|^{\alpha}}{\eta_{+}({a},{0})}\bigwedge\frac{|b_{j}|^{\alpha}}{\eta_{+}({b},{0})}\right).$ (A.58)
That $\lambda(\xi(a),\xi(b))=\lambda(X,Y)$ exists and has the form specified in the statement of the lemma now follows from (A.58) and Lemma A.7, which is applicable because of Lemma A.5. ∎
A.2.10 Proof of Proposition 3.2
Proof.
Recall from (3.28) that
$\lambda(Y_{t+h},\widehat{Y}_{t+h}^{\text{(opt)}})=\sum_{j\geq 0}\kappa_{+}(a_{j+h})\frac{|a_{j+h}|^{\alpha}}{\eta_{+}({a},{0})}.$ (A.59)
On the other hand, by Lemma 3.4,
$\lambda(Y_{t+h},Y_{t})$ $=\sum_{j\geq 0}\kappa_{+,+}(a_{j+h},a_{j})\Big{(}\frac{|a_{j+h}|^{\alpha}}{\eta_{+}({a},{0})}\bigwedge\frac{|a_{j}|^{\alpha}}{\eta_{+}({a},{0})}\Big{)}$ (A.60)
$=\sum_{j\geq 0}\kappa_{+,+}(a_{j+h},a_{j})\frac{|a_{j+h}|^{\alpha}}{\eta_{+}({a},{0})},$ (A.61)
[…]
$=\kappa_{+}(a_{j+h}),$
and hence the right-hand sides of (A.59) and (A.61) are identical. This proves (3.30).
On the other hand, if
$\lambda(Y_{t+h},-Y_{t})$ $=\sum_{j\geq 0}\kappa_{+,-}(a_{j+h},a_{j})\Big{(}\frac{|a_{j+h}|^{\alpha}}{\eta_{+}({a},{0})}\bigwedge\frac{|a_{j}|^{\alpha}}{\eta_{-}({a},{0})}\Big{)},$ (A.62)
For all $j\geq 0$, we have
$\Big{(}\frac{|a_{j+h}|^{\alpha}}{\eta_{+}({a},{0})}\bigwedge\frac{|a_{j}|^{\alpha}}{\eta_{+}({a},{0})}\Big{)}\leq\Big{(}1\bigvee\frac{\eta_{-}({a},{0})}{\eta_{+}({a},{0})}\Big{)}\cdot\Big{(}\frac{|a_{j+h}|^{\alpha}}{\eta_{+}({a},{0})}\bigwedge\frac{|a_{j}|^{\alpha}}{\eta_{-}({a},{0})}\Big{)},$ (A.63)
where we used the fact that $x\wedge(wy)\leq(1\vee w)(x\wedge y)$, for all non-negative $x,y,$ and $w$. At the same time, for the sum of the weights in the right-hand sides of (A.60) and (A.62), we have
$=\kappa_{+}(a_{j+h}),$ (A.64)
[…]
By (3.29), however, we have $|a_{j+h}|^{\alpha}\wedge|a_{j}|^{\alpha}=|a_{j+h}|^{\alpha}$, and in view of (A.59) the left-hand side of the last relation equals $\lambda(Y_{t+h},\widehat{Y}_{t+h}^{\text{(opt)}})$. This completes the proof of the upper bound in (3.31). In view of (A.59), and Relations (A.60), (A.62), and (A.2.10), the lower bound in (3.31) is an immediate consequence of the inequalities:
@@ A.3 Additional Information for Section 4
A.3.1 Extremal Optimal Precisions for FARIMA$(0,d,0)$ Models
[…]
$Y_{t}=(1-B)^{-d}\epsilon_{t},\quad t\in{\mathbb{Z}},$ (A.65)
[…]
For $h=1$ and various pairs of $d$ and $\alpha$ satisfying $\alpha\in(1,2)$, $d\in(0,1-1/\alpha)$, we approximated $\lambda_{h}^{\rm(opt)}$ by truncating the series using their first million terms. The approximations are shown in the left panel of Figure A.2. We see that for a fixed $\alpha$, as $d$ increases, $\lambda_{h}^{\rm(opt)}$ increases. This is intuitive as the level of long-range dependence in $\{Y_{t}\}$ increases as $d$ increases. For example, when $d=0$, (A.65) simplifies to $Y_{t}=\epsilon_{t}$, $t\in{\mathbb{Z}}$, so $\{Y_{t}\}$ cannot be long-range dependent as it consists of independent random variables. The higher the level of long-range dependence, the more $Y_{t+h}$ depends on $Y_{s}$’s in the distant past, and the better the optimal predictor performs. On the other hand, for a fixed $d$, $\lambda_{h}^{\rm(opt)}$ decreases as $\alpha$ increases. The right panel of Figure A.2 shows that for $\alpha=1.4$ and $d=0.19$, $\lambda_{h}^{\rm(opt)}$ decreases as the forecast horizon increases, as one would expect.
Figure A.2: Left: a contour plot showing how the one-step-ahead extremal optimal precision for the FARIMA(0, $d$, 0) model with symmetric $\alpha$-stable innovations (A.65) varies as $\alpha$ and $d$ vary. Right: the extremal optimal precision as a function of $h$, for $(\alpha,d)=(1.4,0.19)$, which is represented by the red dot in the left panel.
A.3.2 Further Data Analysis Results
@@
$h$ $p$ Baseline FARIMA OLS AR LAD AR Baseline FARIMA OLS AR LAD AR
1 0.90 0.492 0.448 0.354 0.442 0.435 0.392 0.276 0.419
1 0.95 0.379 0.368 0.258 0.332 0.334 0.320 0.225 0.352
1 0.99 0.244 0.238 0.104 0.148 0.262 0.262 0.124 0.230
6 0.90 0.308 0.339 0.284 0.302 0.256 0.298 0.215 0.293
6 0.95 0.189 0.213 0.150 0.157 0.173 0.198 0.138 0.174
6 0.99 0.049 0.119 0.125 0.076 0.040 0.115 0.162 0.131
12 0.90 0.284 0.292 0.253 0.282 0.197 0.205 0.152 0.216
12 0.95 0.194 0.176 0.127 0.181 0.148 0.124 0.076 0.176
12 0.99 0.098 0.073 0.049 0.028 0.085 0.061 0.056 0.029
18 0.90 0.270 0.309 0.253 0.279 0.210 0.256 0.183 0.260
18 0.95 0.175 0.197 0.142 0.165 0.154 0.173 0.117 0.184
18 0.99 0.049 0.116 0.067 0.060 0.040 0.115 0.064 0.083
Table A.3: An expanded summary of the performance of various predictors. The order of the AR models was 168. For each pair of $h$ and $p$, the goal was to predict exceedance of the $p$th marginal quantile $h$ steps ahead. In each row, for each metric, the largest value is bolded. The LAD numbers are bigger than the corresponding OLS numbers, with the exception of the LAD precisions for $h\in\{6,12,18\},p=0.99$ and the LAD TSS values for $h\in\{6,12\},p=0.99$. We believe that the LAD predictor performed better in most cases because it had a higher alarm rate in most cases; see Figure A.3.
Figure A.3: A comparison of the test set alarm rates for the AR models fit using OLS and LAD. See Table A.3 for their precisions and TSS values. Different points represent different lead times $h$. Given $p$, the test alarm rate for the $p$th quantile is defined as the proportion of times in the test set for which an exceedance of the $p$th quantile was predicted. Along the dashed line, the OLS and LAD rates are equal. For every pair of $h$ and $p$, the LAD rate is higher than the OLS rate. A higher alarm rate corresponds to a lower threshold for predicting extremeness, which yields a higher TPR and a higher FPR. In this case, the changes led to increases in the precision and TSS for most pairs of $h$ and $p$.